# On the Algorithmic Bias of Aligning Large Language Models with RLHF: Preference Collapse and Matching Regularization

### v1

@@ 3.2 Preference Matching Regularization
In this case, (3.9) reduces to
$\max_{\phi}\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}\big{[}r(x,y)-\log(\pi_{\phi}(y|x))\big{]}.$ (3.10)
@@
We can interpret the PM RLHF objective (3.10) through the lens of robust optimization. Note that the dual of (3.10) is
@@
where $\text{Unif}(y|x))$ denotes the uniform distribution. We call this uniform-penalty RLHF, which can be thought of as an instance of the standard KL RLHF in (2.3) when $\beta=1$ and $\pi_{\textnormal{ref}}(y|x)$ is the uniform distribution.
@@
3.3 Extension to Response-Dependent Regularization
This, in particular, includes the standard KL RLHF (2.3).
Just as with its response-independent counterpart, Hypothesis 3.1 allows us to write (3.11) in the following $\phi$-independent form:
$\max_{\pi}\mathbb{E}_{y\sim\pi(\cdot|x)}\big{[}r(x,y)+R_{x,y}(\pi(y|x))\big{]}.$
Next, Hypothesis 3.2 shows that the $\phi$-independent optimization program above can be further reduced to
To admit the PM policy as its solution, just as with (3.7), the optimization program (3.12) shows that $R_{i}$ must obey
$R_{i}(\pi)+\pi R_{i}^{\prime}(\pi)+\log\pi=c^{\prime}$
for all $1\leq i\leq k$. Solving this differential equation gives
where $C_{1}=c^{\prime}-1$ is independent of $i$, while $C_{2,i}$ could depend on $i$. As the derivation of (3.13) is conditional on $x$, both the constants $C_{1}$ and $C_{2,i}$ can depend on $x$. Taken together, our discussion above suggests the following theorem.
Theorem 2.
$R_{x,y}(\pi)=-\log\pi+C_{1,x}+\frac{C_{2,x,y}}{\pi},$
The sufficient part of the proof of this theorem is almost the same as that of Theorem 1. It is thus omitted.
Remark 3.6.
For the last term in $R_{x,y}(\pi)$, note that
$\mathbb{E}_{y\sim\pi(\cdot|x)}\frac{C_{2,x,y}}{\pi(y|x)}=\mathbb{E}_{x}\left[\sum_{y}\pi(y|x)\frac{C_{2,x,y}}{\pi(y|x)}\right]=\mathbb{E}_{x}\left[\sum_{i=1}^{k}C_{2,x,y_{i}}\right]$
is a constant that is independent of the decision variable $\pi$.
Proposition 3.7.
The standard KL RLHF (2.3) does not yield the PM policy.
3.4 Extension to Generalized Preference Models
From a statistical standpoint, the use of the exponential function in the BTL and PL models is by no means the only choice for modeling preference probabilities. In light of this, it would be interesting to develop PM RLHF with generalized reward modeling. In particular, we consider a generalized BTL model where the human preference is driven by an implicit reward model $r(x,y)$, but the exponential function is replaced by a positive and increasing function $g$. Then, the preference probability in the generalized BTL model becomes
$\mathbb{P}(y_{1}\succ y_{2}|y_{1},y_{2},x)=\frac{g(r(x,y_{1}))}{g(r(x,y_{1}))+g(r(x,y_{2}))}.$
Accordingly, the generalized PL model becomes
$\mathbb{P}\big{(}y|y_{1},y_{2},\ldots,y_{k},x\big{)}=\frac{g(r(x,y))}{g(r(x,y_{1}))+g(r(x,y_{2}))+\cdots+g(r(x,y_{k}))}.$
For simplicity, we consider response-independent regularization as in Section 3.2. Assuming differentiability of $g$, in the appendix, we obtain the PM differential equation for the generalized PL model:
$\pi R^{\prime\prime}(\pi)+2R^{\prime}(\pi)+\frac{1}{g^{\prime}(g^{-1}(\pi))}=0.$
Our PM regularization can be extended by the following result, the proof of which is deferred to the appendix.
Theorem 3.
Under the same assumptions as in Theorem 1, consider using the generalized PL model for modeling the preference probabilities in (3.1). Denoting $G(r)=\int g(r)\mathrm{d}r$, $R$ is a PM regularizer if and only if
$R(\pi)=-g^{-1}(\pi)+C_{1,x}+\frac{G(g^{-1}(\pi))+C_{2,x}}{\pi},$ (3.14)
where $C_{1,x}$ and $C_{2,x}$ are arbitrary constants depending on $x$.
Remark 3.8.
The indefinite integral $\int g(r)\mathrm{d}r$, by definition, is unique only up to a constant. This is reflected by the constant $C_{2,x}$ in (3.14). In the case of the original PL model, $g(r)=\mathrm{e}^{r}$, the additional term $G(g^{-1}(\pi))$ becomes $\exp(\log\pi)/\pi=\pi/\pi=1$. As such, the numerator $G(g^{-1}(\pi))+C_{2,x}$ in (3.14) becomes a constant, which is consistent with Theorem 1.
If a generalized PL model can better capture human preferences, then this PM regularization is likely to enhance the performance of preference alignment. As for what function $g$ better captures human preference, we leave this question for future investigation.
3.5 Proof of Theorem 1
The discussion before the statement of Theorem 1 shows the necessity of (1). We finish the proof below by showing the sufficiency of (1). The objective function is:
$\mathbb{E}_{y\sim\pi(\cdot|x)}r(x,y)-\log(\pi(y|x)).$
$=$ $-D_{\text{KL}}(\pi(y|x)\|\pi^{\star}(y|x))+\mathbb{E}_{x}\log Z(x).$
It is also independent of $\pi(y|x)$. Therefore, we obtain
4 Conditional Preference Matching RLHF
When applying PM RLHF in practice, we encounter a numerical issue that is inherent to the nature of text data. A direct application of PM RLHF in our experiments fails to produce human-like text. For example, Table 4 in the appendix presents two examples of responses where the LLM struggles to produce sentences that read naturally. This numerical issue corresponds to high values of perplexity [^8] Letting $\pi(y|x)=p_{1}\cdot p_{2}\cdots p_{\text{Length}(y)}$, where $p_{i}$ is the probability of generating the $i^{\textnormal{th}}$ token given the previous tokens, $\text{Perplexity}=[p_{1}\cdot p_{2}\cdots p_{\text{Length}}(y)]^{-1/\text{Length}(y)}$. A small value generally implies good quality of the text. A formal description is given in Section 6., as illustrated by the experimental results in Figure 1. In particular, we evaluate the perplexity of the PM RLHF–aligned and KL RLHF–aligned models on 12,500 test samples. For the OPT-1.3B model (Zhang et al., 2022) fine-tuned by PM RLHF, the perplexity of the LLM is 2343.56, while KL RLHF yields a significantly lower perplexity of 7.24. Similarly, for the Llama-2-7B model (Touvron et al., 2023), the perplexity scores for LLMs trained by PM RLHF and KL RLHF are 7.15 and 5.48, respectively.
A possible cause is that the number of all responses is excessively large. In principle, the number can be up to $|\mathcal{V}|^{L}$ for responses of length no more than $L$, where $\mathcal{V}$ is the token vocabulary. For example, $|\mathcal{V}|=50,272$ for the OPT-1.3B model, and $|\mathcal{V}|=32,000$ for the LLaMA series models, and $|\mathcal{V}|=50,257$ for GPT-2 and GPT-3.5 series models (OpenAI, 2023). However, most such responses are meaningless and the meaningful and natural text occupies only a negligible fraction. In particular, the reward model might not be generalized well on the meaningless text. Consequently, PM RLHF does not exclude the generation of meaningless text as the reward model might assign positive probability to such text.
4.1 Conditional Preference Matching
To tackle the numerical issue above, we propose to divide the responses space, which we refer to as $\mathcal{Y}$, into $\mathcal{M}(x)$ and $\mathcal{Y}\setminus\mathcal{M}(x)$. The former contains all meaningful responses given $x$, and the latter contains all meaningless responses given $x$. While the form of these sets is to be determined, we assume for/at the moment that the reward model behaves well on $\mathcal{M}(x)$ in the sense that the rewards on this set reflects well the human labelers’ preferences.
Our hypothesis is that by conditioning PM RLHF on the set where the reward model behaves well, the numerical issue would be resolved. As such, we consider a constraint variant of (3.9): [^9] The constraint $\pi(y|x)\geq 0$ is automatically satisfied by parametrization of neural networks.
s.t. $\pi(y|x)\geq 0,\ \text{if}\ y\in\mathcal{M}(x),$
$\pi(y|x)=0,\ \text{if}\ y\notin\mathcal{M}(x).$
However, this program is difficult to optimize due to a large number of constraints. Instead, we consider to an unconstraint relaxation of the above optimization program:
where $\epsilon>0$ is a small number.
The following theorem shows that the global solution to (4.1) is preference matching in the conditional sense.
Theorem 4.
$\pi^{\star}(y|\mathcal{M}(x),x)=\frac{\mathrm{e}^{r(x,y)}}{\sum_{y^{\prime}\in\mathcal{M}(x)}\mathrm{e}^{r(x,y^{\prime})}}$
for any $y\in\mathcal{M}(x)$.
This theorem holds for an arbitrary choice of $\mathcal{M}(x)$, not necessarily that containing only meaningful responses for the prompt $x$. In particular, by letting $\epsilon\rightarrow 0$, $\pi^{\star}(y|x)$ tend to output responses only in the meaningful set $\mathcal{M}(x)$, that is, $\pi^{\star}(\mathcal{M}(x)|x)\rightarrow 1$. As such, we have
$\pi^{\star}(y|x)=\frac{\mathrm{e}^{r(x,y)}}{\sum_{y^{\prime}\in\mathcal{M}(x)}\mathrm{e}^{r(x,y^{\prime})}}$
in the limit $\epsilon\rightarrow 0$.
4.2 Parameterization of Conditional PM RLHF
To conclude this section, we discuss some practical aspects of implementing the conditional PM RLHF. Following Example 2 in Section 3.2, we set $C_{1,x}=C_{1}=-\mathbb{E}_{y\sim\pi(\cdot|x)}r(x,y)$, where the expectation is over the randomness of both $x$ and $y$, and $C_{2,x}=0$ (Ouyang et al., 2022). The reward now has mean zero and is still denoted as/by $r(x,y)$ for simplicity. The objective function becomes
To implement the conditional PM RLHF above, both the regular set $\mathcal{M}(x)$ and the relaxation parameter $\epsilon$ need to be set. We leverage the reference model to serve as a basis to characterize regular responses. Specifically, we assume that a regular response would yield a relatively large value of $\pi_{\textnormal{ref}}(y|x)$, which suggests setting the regular set as $\mathcal{M}(x):=\{y:\pi_{\textnormal{ref}}(y|x)\geq\alpha\}$ for some $\alpha>0$. It is worthwhile mentioning that Theorem 4 remains true/correct for this choice of regular set.
Next, we discuss how to choose $\epsilon$. First of all, while its choice might influence the likelihood of generating a response, it does not impact the conditional preference matching of those responses such that $\pi_{\textnormal{ref}}(y|x)\geq\alpha$. In general, a very small value of $\epsilon$ is desired from a theoretical perspective since it penalizes the generation of meaningless responses. However, this meanwhile increases the burden of numerical optimization. To balance, we let $\epsilon$ depend on both $x,y$ and set $\epsilon(x,y)=\pi_{\textnormal{ref}}(y|x)$ to serve as a computationally feasible calibration for relaxation. This would ensure that the model remains reasonably close to the reference model, avoiding significant deviations that could impact the model robustness and reliability. Finally, we propose the following conditional PM RLHF that we will use in our experiments:
Definition 4.1.
We refer to
as the conditional PM RLHF.
To conclude this section, we compare several variants of RLHF, including ours, in Table 1.
Table 1: Comparison of RLHF with different regularizers. $f$-divergence RLHF will be introduced in Section 5. Notably, Maximum Entropy RLHF is generally not preference matching unless for a certain choice of its tuning parameter. The proposed conditional PM RLHF is the only RLHF objective that simultaneously achieves natural language generation and preference matching.
Natural Language Generation Preference Matching
Standard (KL) RLHF
$f$-divergence RLHF
Maximum Entropy RLHF ✗/✓
PM RLHF (Ours)
Conditional PM RLHF (Ours)
5 Algorithmic Biases of Standard RLHF
5.1 Bias from KL Regularization
Proposition 5.1 (Rafailov et al. (2023)).
The optimal solution of the RLHF problem to (2.3) is
$\pi_{\textnormal{rlhf}}(y|x)=\frac{\pi_{\textnormal{ref}}(y|x)\mathrm{e}^{r(x,y)/\beta}}{\sum_{j}\pi_{\textnormal{ref}}(y_{j}|x)\mathrm{e}^{r(x,y_{j})/\beta}}.$ (5.1)
The closed-form solution provided in (5.1) has been previously presented in the literature (Rafailov et al., 2023). However, we argue that (5.1) inherently leads to a certain algorithmic bias. To show this bias, it is instructive to consider the binary setting of (5.1). Given a prompt $x$ and two responses $y_{1}$ and $y_{2}$, we denote
Similarly, we denote by
$p_{\textnormal{reward}}(y_{i})=\frac{\mathrm{e}^{r(x,y_{i})/\beta}}{\mathrm{e}^{r(x,y_{1})/\beta}+\mathrm{e}^{r(x,y_{2})/\beta}},\ p_{\textnormal{ref}}(y_{i})=\frac{\pi_{\textnormal{ref}}(y_{i}|x)}{\pi_{\textnormal{ref}}(y_{1}|x)+\pi_{\textnormal{ref}}(y_{2}|x)},\ i=1,2.$
Here all the probabilities $p_{\text{rlhf}}(y_{i})$, $p_{\textnormal{reward}}(y_{i})$ and $p_{\textnormal{ref}}(y_{i})$ are conditioned on $x,y_{1},y_{2}$. For the sake of simplifying the notation and improving readability, we omit the conditioning variables. Then, based on (5.1), the probability of binary comparison of the RLHF model is
$p_{\text{rlhf}}(y_{i})=$ $\frac{\pi_{\textnormal{rlhf}}(y_{i}|x)}{\pi_{\textnormal{rlhf}}(y_{1}|x)+\pi_{\textnormal{rlhf}}(y_{2}|x)}$ (5.2)
$=$ $\frac{\pi_{\textnormal{ref}}(y_{i}|x)\mathrm{e}^{r(x,y_{i})/\beta}}{\pi_{\textnormal{ref}}(y_{1}|x)\mathrm{e}^{r(x,y_{1})/\beta}+\pi_{\textnormal{ref}}(y_{2}|x)\mathrm{e}^{r(x,y_{2})/\beta}}$
$=$ $\frac{p_{\textnormal{ref}}(y_{i})p_{\textnormal{reward}}(y_{i})}{p_{\textnormal{ref}}(y_{1})p_{\textnormal{reward}}(y_{1})+p_{\textnormal{ref}}(y_{2})p_{\textnormal{reward}}(y_{2})},\ i=1,2.$
From (5.2), it is clear that $p_{\text{rlhf}}$ is identical to the preferences of the reward model when $\beta=1$ and, more crucially, $\pi_{\textnormal{ref}}$ is the uniform distribution over all (valid) responses for any prompt $x$. Figure 2 illustrates this algorithmic bias by showing the relationship between the reward model and the RLHF policy $\pi_{\textnormal{rlhf}}(y|x)$ in terms of their preference probabilities. As long as the reference model is not uniform between the two responses—that is, $p_{\textnormal{ref}}(y_{1})\neq p_{\textnormal{ref}}(y_{2})\neq 0.5$—the two models are not aligned. In extreme case when $p_{\textnormal{ref}}(y_{1})=0$ or 1, KL RLHF will lead to a phenomenon we refer to as preference collapse and will be discussed in Section 5.2.
As mentioned in Example 3 in Section 3.2, however, it is practically impossible to have a uniform distribution for a reference language model. Specifically, $\pi_{\textnormal{ref}}(y|x)$ represents a pretrained model or SFT model, which by no means yields a uniform distribution. Early stopping can reduce this algorithmic bias to some extent (Schulman, 2023) but cannot completely get rid of it as the target of (5.1) is fundamentally biased. In stark contrast, conditional PM RLHF provably sets its global solution to be unbiased with respect to the reward model.
This algorithmic bias stems from the use of a reference model in the RL objective (5.1) and, in particular, cannot be removed by replacing the KL divergence with other divergences. To see this point, consider the $f$-divergence RLHF (Wang et al., 2023):
@@ 5.1 Bias from KL Regularization
$p^{f}_{\textnormal{rlhf}}(y_{i})=\frac{p_{\textnormal{ref}}(y_{i})p^{f}_{\textnormal{reward}}(y_{i})}{p_{\textnormal{ref}}(y_{1})p^{f}_{\textnormal{reward}}(y_{1})+p_{\textnormal{ref}}(y_{2})p^{f}_{\textnormal{reward}}(y_{2})},\ i=1,2,$ (5.3)
@@
As is clear from (5.3), the algorithmic bias remains in the case of $f$-divergence. In fact, it becomes worse in the sense that the bias remains even if $\beta=1$ and $p_{\textnormal{ref}}$ is uniform, unless $(f^{\prime})^{-1}(u)=\exp(u+c)$ for some constant $c$. Note that $(f^{\prime})^{-1}(u)=\exp(u+c)$ if and only if $f(u)=u\log u+c^{\prime}u$, which effectively gives the KL divergence.
5.2 Extreme Case: Preference Collapse in KL RLHF
@@
This extreme case is also shown in Figure 2. The phenomenon of preference collapse, if were true, would be detrimental to decision-making as it would completely let the LLM discard minority opinions. Our experiments on Llama-2-7B and OPT-1.3B indeed show this phenomenon occurs in practice, with detailed experimental setup given in Section 6. For each prompt $x$ in the dataset, we randomly generate two responses, $y_{1}$ and $y_{2}$. We then calculate the conditional probabilities of $y_{1}$ with respect to $y_{1}$ and $y_{2}$. As shown in Figure 3(a), it becomes evident that the probabilities $p_{\textnormal{ref}}(y_{1})$ are very likely to be close to 0 or 1. In contrast, the probabilities induced by the reward model are more uniform, with probabilities between 0 and 1 for a large number of prompts, as illustrated in Figure 3(b). This phenomenon can be observed in the experiment on OPT-1.3B, as demonstrated in Figure 4.
While the exact cause is unclear to us at the moment, some evidence shows that it might be attributed to varying lengths of responses (Park et al., 2024; Chen et al., 2024a). Recall that LLMs generate each token following the next-token prediction. The probability of a response is in the form of the product of probabilities and the number of terms is the length of the response. A direct consequence is that the probability of a response in general decays exponentially with its length. Thus, for two responses that are generated independently from a prompt, the ratio of their probability is likely to be either very large or small, which is especially the case when the two responses differ significantly in length.

### v2

@@ A.4 Examples of Preference Matching Regularization
In this case, (A.3) reduces to
$\max_{\phi}\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}\big{[}r(x,y)-\log(\pi_{\phi}(y|x))\big{]}.$ (A.4)
@@
We can interpret the PM RLHF objective (A.4) through the lens of robust optimization. Note that the dual of (A.4) is
@@
where $\text{Unif}(y|x))$ denotes the uniform distribution. We call this uniform-penalty RLHF, which can be thought of as an instance of the standard KL RLHF in (A.2) when $\beta=1$ and $\pi_{\textnormal{ref}}(y|x)$ is the uniform distribution.
@@
A.5 Discussion of $f$-divergence RLHF
The algorithmic bias stems from the use of a reference model in the RL objective (5.1) and, in particular, cannot be removed by replacing the KL divergence with other divergences. To see this point, consider the $f$-divergence RLHF (Wang et al., 2023):
@@ A.5 Discussion of $f$-divergence RLHF
$p^{f}_{\textnormal{rlhf}}(y_{i})=\frac{p_{\textnormal{ref}}(y_{i})p^{f}_{\textnormal{reward}}(y_{i})}{p_{\textnormal{ref}}(y_{1})p^{f}_{\textnormal{reward}}(y_{1})+p_{\textnormal{ref}}(y_{2})p^{f}_{\textnormal{reward}}(y_{2})},\ i=1,2,$ (A.5)
@@
As is clear from (A.5), the algorithmic bias remains in the case of $f$-divergence. In fact, it becomes worse in the sense that the bias remains even if $\beta=1$ and $p_{\textnormal{ref}}$ is uniform, unless $(f^{\prime})^{-1}(u)=\exp(u+c)$ for some constant $c$. Note that $(f^{\prime})^{-1}(u)=\exp(u+c)$ if and only if $f(u)=u\log u+c^{\prime}u$, which effectively gives the KL divergence.
A.6 Other Discussions
Generating Examples for PM RLHF.
Human: What was the name of the movie that won the Academy Award for Best Picture in
1955?
Assistant: The movie that
Human: Why do birds migrate south for the winter?
Assistant: The birds migrate south for the winter because the climate in the northern part of
the world is the most suitable for the birds to live in. The climate in the northern part of the
world is the most suitable for the birds to live in.
Human: thank thank thank thank thank thank thank thank thank thank thank thank thank
thank thank thank thank thank thank thank thank thank thank thank thank thank thank
thank thank thank
Abbreviation Conditional Probability Formulation (i=$1,2$)
$p_{\text{rlhf}}(y_{i})$ $p_{\text{rlhf}}(y_{i}|y_{1},y_{2},x)$ $\pi_{\text{rlhf}}(y_{i}|x)/(\pi_{\text{rlhf}}(y_{1}|x)+\pi_{\text{rlhf}}(y_{2}|x))$
$p_{\textnormal{ref}}(y_{i})$ $p_{\textnormal{ref}}(y_{i}|y_{1},y_{2},x)$ $\pi_{\textnormal{ref}}(y_{i}|x)/(\pi_{\textnormal{ref}}(y_{1}|x)+\pi_{\textnormal{ref}}(y_{2}|x))$
$p_{\textnormal{reward}}(y_{i})$ $p_{\textnormal{reward}}(y_{i}|y_{1},y_{2},x)$ $\exp(r(x,y_{i})/\beta)/(\exp(r(x,y_{1})/\beta)+\exp(r(x,y_{2})/\beta))$
Appendix B Algorithmic Biases of Standard RLHF
B.1 Bias from KL Regularization
Proposition B.1 (Rafailov et al. (2023)).
The optimal solution of the RLHF problem to (2.3) is
$\pi_{\textnormal{rlhf}}(y|x)=\frac{\pi_{\textnormal{ref}}(y|x)\exp(r(x,y)/\beta)}{\sum_{j}\pi_{\textnormal{ref}}(y_{j}|x)\exp(r(x,y_{j})/\beta)}.$ (B.1)
The closed-form solution provided in (B.1) has been previously presented in the literature of direct preference optimization (Rafailov et al., 2023). However, we argue that (B.1) inherently leads to a certain algorithmic bias. To show this bias, it is instructive to consider the binary setting of (B.1). Given a prompt $x$ and two responses $y_{1}$ and $y_{2}$, we denote
Similarly, we denote by
$p_{\textnormal{reward}}(y_{i})=\frac{\exp(r(x,y_{i})/\beta)}{\exp(r(x,y_{1})/\beta)+\exp(r(x,y_{2})/\beta)},\ p_{\textnormal{ref}}(y_{i})=\frac{\pi_{\textnormal{ref}}(y_{i}|x)}{\pi_{\textnormal{ref}}(y_{1}|x)+\pi_{\textnormal{ref}}(y_{2}|x)},\ i=1,2.$
Here, all the probabilities $p_{\text{rlhf}}(y_{i})$, $p_{\textnormal{reward}}(y_{i})$, and $p_{\textnormal{ref}}(y_{i})$ are conditioned on $x,y_{1},y_{2}$. For the sake of simplifying the notation and improving readability, we omit the conditioning variables. Then, based on (B.1), the probability of binary comparison of the RLHF model is
$p_{\text{rlhf}}(y_{i})=\frac{p_{\textnormal{ref}}(y_{i})p_{\textnormal{reward}}(y_{i})}{p_{\textnormal{ref}}(y_{1})p_{\textnormal{reward}}(y_{1})+p_{\textnormal{ref}}(y_{2})p_{\textnormal{reward}}(y_{2})},\ i=1,2.$ (B.2)
From (B.2), it is clear that $p_{\text{rlhf}}$ is identical to the preferences of the reward model when $\beta=1$ and, more crucially, when $\pi_{\textnormal{ref}}$ is the uniform distribution over all (valid) responses for any prompt $x$. Figure 4 illustrates this algorithmic bias by showing the relationship between the reward model and the RLHF policy $\pi_{\textnormal{rlhf}}(y|x)$ in terms of their preference probabilities. As long as the reference model is not uniform between the two responses—that is, $p_{\textnormal{ref}}(y_{1})\neq p_{\textnormal{ref}}(y_{2})\neq 0.5$—the two models are not aligned. In an extreme case when $p_{\textnormal{ref}}(y_{1})=0$ or 1, KL RLHF will lead to a phenomenon—that, as mentioned earlier—we refer to as preference collapse and will be discussed in Section B.2.
As mentioned in Example 3 later in Section A.4, however, it is practically impossible to have a uniform distribution for a reference language model. Specifically, $\pi_{\textnormal{ref}}(y|x)$ represents a pretrained model or SFT model, which by no means yields a uniform distribution. Early stopping can reduce this algorithmic bias to some extent (Schulman, 2023) but cannot completely remove it, as the target of (B.1) is fundamentally biased. In stark contrast, conditional PM RLHF provably sets its global solution to be unbiased with respect to the reward model.
B.2 Extreme Case: Preference Collapse in KL RLHF
@@
This extreme case is also shown in Figure 4. The phenomenon of preference collapse, if were true, would be detrimental to decision-making, as it would completely let the LLM discard minority opinions. Our experiments on Llama-2-7B and OPT-1.3B indeed show this phenomenon occurs in practice. For each prompt $x$ in the dataset, we randomly generate two responses, $y_{1}$ and $y_{2}$. We then calculate the conditional probabilities of $y_{1}$ with respect to $y_{1}$ and $y_{2}$. As shown in Figure 5(a), it becomes evident that the probabilities $p_{\textnormal{ref}}(y_{1})$ are very likely to be close to 0 or 1. In contrast, the probabilities induced by the reward model are more uniform, with probabilities between 0 and 1 for a large number of prompts, as illustrated in Figure 5(b). This phenomenon can be observed in the experiment on OPT-1.3B, as demonstrated in the appendix.
While the exact cause of this phenomenon is unclear to us at the moment, some evidence shows that it might be attributed to varying lengths of responses (Park et al., 2024; Chen et al., 2024a). Recall that LLMs generate each token following the next-token prediction. The probability of a response is in the form of the product of probabilities and the number of terms is the length of the response. A direct consequence is that the probability of a response in general decays exponentially with its length. Thus, for two responses that are generated independently from a prompt, the ratio of their probability is likely to be either very large or small, which is especially the case when the two responses differ significantly in length.